# Little By Little: Continual Learning via Self-Activated Sparse Mixture-of-Rank Adaptive Learning

### v1

@@ Abstract
Continual learning (CL) with large pre-trained models is challenged by catastrophic forgetting and task interference. Existing LoRA-based Mixture-of-Experts (MoE) approaches mitigate forgetting by assigning and freezing task-specific adapters, but suffer from interference, redundancy, and ambiguous routing due to coarse adapter-level selection. However, this design introduces three key challenges: 1) Interference: Activating full LoRA experts per input leads to subspace interference and prevents selective reuse of useful components across tasks. 2) Redundancy: Newly added experts often duplicate or contradict existing knowledge due to unnecessary activation of unrelated ranks and insufficient reuse of relevant ones. 3) Ambiguity: Overlapping features across tasks confuse the router, resulting in unstable expert assignments. As more experts accumulate, earlier task routing degrades, accelerating forgetting. We propose MoRA, a Mixture-of-Rank Adaptive learning approach with self-activated and sparse rank activation for CL. Unlike mixing multiple low-rank matrices, MoRA decomposes each rank-r update into r rank-1 components, each treated as an independent expert, enabling fine-grained mixture of rank-1 expert utilization while mitigating interference and redundancy. To avoid ambiguous routing, we propose that each rank-1 expert can infer its own relevance via intermediate activations. Coupled with our proposed rank pruning and activation budgets, MoRA adaptively selects a sparse mixture of ranks per input. We validate MoRA on continual learning tasks with CLIP and large language models (LLMs), analyzing both in-domain learning and out-of-domain forgetting/generalization during fine-tuning. MoRA shows significant effectiveness on enhancing CL with PTMs, and improving generalization while mitigating forgetting.
@@ 1 Introduction
Continual learning (CL) [29, 17, 73, 43, 21] aims to enable models to incrementally and efficiently acquire new knowledge from a stream of tasks, without catastrophic forgetting [55, 51] or the need for repeated fine-tuning on all previously seen data [77, 76, 85, 72]. The emergence of large pre-trained models (PTMs), including Vision Transformer (ViT) for vision [22], Large Language Model (LLM) [28, 61], and CLIP for vision–language embeddings [60, 33], has spurred a wave exploring CL with PTMs[94, 52, 72, 35, 81, 59, 62, 74, 58]. Despite the success of PTMs, CL on them across sequential tasks still faces key challenges, including preventing catastrophic forgetting and preserving—or even enhancing—the pre-trained capabilities to generalize to unseen or future cases [92, 85, 48, 58].
@@
A common approach to integrating new knowledge into pre-trained models (PTMs) is continual full fine-tuning [92, 27] or parameter-efficient fine-tuning (PEFT) using a small set of learnable parameters, such as Low-Rank Adaptation (LoRA) [32] (Fig. 1(a)). However, continually updating the same set of parameters often causes interference and forgetting, even with a single new task [7]. Leveraging LoRA’s inherent regularization and flexibility, many CL methods for PTMs employ LoRA or its variants, along with specialized strategies to mitigate forgetting [72, 44, 74, 58, 44, 74].
@@
Inspired by mixture-of-experts (MoE) architectures [65, 40, 24, 16] and the modular nature of LoRA adapters, MoE frameworks using LoRA as experts (Fig. 1(b)) have been widely adopted in CL [85, 72, 84, 13]. These methods incrementally train new LoRA experts while freezing previous ones, using a router to compose experts per sample. This isolation reduces cross-task interference and forgetting, while the composability of experts supports knowledge reuse and transfer [72, 58, 64]. In MoE-LoRA models, each expert comprises a full LoRA adapter with multiple ranks (i.e., subspaces), and expert selection occurs at this coarse-grained adapter level. Despite their success, this coarse granularity and the routing mechanism introduce three key challenges that limit effective utilization:
(1) Interference across experts and infeasible knowledge reuse. For any given sample, an entire LoRA expert must be either fully activated or not at all, leading to two major issues: interference among incompatible ranks from different experts, and limited flexibility in knowledge reuse.
@@
(2) Learning redundancy and conflicts. Newly added experts may relearn redundant information or learn conflicting representation due to unnecessary activation of unrelated old ranks or insufficient reuse of relevant ones. This coarse selection exacerbates interference and forgetting during inference.
(3) Routing ambiguity and forgetting. Redundancy and conflicts among experts make it hard for a LoRA-level router to assign overlapping or common inputs correctly, leading to suboptimal mixtures [40, 24, 16]. As more new experts are accumulated in CL, the router’s assignments for earlier tasks become increasingly unreliable, makes it more prone to forgetting.
@@
We propose MoRA, a Mixture-of-Rank Adaptive learning approaches with self-activated and sparsely selected ranks across the LoRAs added at different tasks. To address the limitations of coarse-grained MoE-LoRA methods, MoRA operates at a fine-grained level by treating each rank-one subspace within a LoRA adapter as an individual expert. Expert selection and routing are performed at the rank level, reducing interference from activating entire LoRA blocks. This decomposition promotes specialization on narrow regions of the data manifold, reduces redundancy, and enables effective reuse of shared knowledge across tasks.
We introduce a self-activation mechanism to select a sparse set of ranks without relying on a separately trained router, avoiding routing confusion and further mitigating forgetting (Fig. 1(c)). Each rank evaluates its own relevance to the input and determines its contribution, grounded in the view of low-rank updates as linear key–value memory. Sparsity is enforced via temperature-scaled softmax together with top-k masking and rank pruning under a rank activation budget. This design introduces no additional computation and eliminates the overhead associated with extra routing. Leveraging the connection between PTMs and low-dimensional manifolds [41, 1], MoRA enables precise, input-specific adaptation at test time through selective subspace updates. With fine-grained rank-level experts and sparse activation, MoRA facilitates continual learning by reusing relevant old ranks and selectively activating a few new ones per sample, enabling efficient, “little-by-little” adaptation that minimizes redundancy and supports essential updates. This selective, sample-specific activation at inference also enhances retention of pre-trained knowledge and improves generalization to unseen or future cases.
We apply and validate MoRA on CL with both vision-language CLIP and LLMs, as well as analyzing the effects on general pre-trained knowledge during low-rank fine-tuning. Across all settings, MoRA achieves strong downstream performance with significantly fewer active parameters, substantially reduces forgetting of both prior tasks and pre-trained knowledge, and even enhances generalization to unseen domains.
@@ 2 Related Work
Continual learning enables sequential knowledge acquisition without forgetting. Experience replay (ER) methods [49, 4, 10, 45, 11, 82, 83] interleave past examples with new data. Parameter regularization [36, 2, 86, 3, 34] penalizes updates to critical weights. Dynamic networks [70, 71, 93, 72, 52, 44, 77, 76, 66, 75, 72] allocate new capacity on the fly and preserve dedicated pathways for prior tasks.
Continual learning of PTMs. For CL on vision–language CLIP model [35, 89, 27], methods like ZSCL [92] retain zero-shot performance during adaptation, and follow-up work [85, 67, 81, 48] continually fine-tunes while leveraging frozen pre-trained predictions. The X-TAIL benchmark [81] further challenges models by mixing domain labels at test time. In large language models (LLMs) [18, 59, 62, 74, 58], continual learning uses capacity expansion or task-specific submodules to reduce interference. In this work, we also study how fine-tuning impacts pre-trained general knowledge.
Low-rank adaptation (LoRA) [32] is widely used for parameter-efficient fine-tuning of large pre-trained models. Building on this foundation, recent methods reformulate LoRA’s updates via SVD-based initialization and dynamic rank scheduling [87, 53, 20, 88, 46, 79], showing that task adaptation primarily lies in a compact subspace. In this work, we offer a complementary perspective by viewing both pre-trained weight matrix and its low-rank updates as a linear associative memory [41, 1, 37, 5]. Under this lens, a rank-$r$ update corresponds to writing $r$ new key–value entries into the pre-trained memory matrix, with each rank-1 component acting as an independent memory slot.
Mixture-of-experts (MoE) in LoRA fine-tuning. MoE increases model capacity by replacing the Transformer’s feed-forward layer with expert subnetworks, routing each input to a small subset and using load-balancing losses to even out usage [65, 40, 24, 16]. This paradigm has been adapted for standard fine-tuning [14, 42, 23] and CL [85, 72, 84, 13] with LoRA, treating each low-rank adapter as an expert: new adapters are trained and frozen per task to prevent forgetting. In contrast, we decompose each rank-$r$ update into $r$ rank-$1$ components and compute an input-dependent mixture over these fine-grained experts, greatly enhancing both expert specialization and mixture diversity.
@@ 3.1 Preliminaries
Continual learning. In continual learning, a model must sequentially learn from $T$ distinct tasks. For task $t\in\{1,\dots,T\}$, we denote its dataset by $\mathcal{D}^{t}=\{(\mathbf{x}^{t}_{i},y^{t}_{i})\}_{i=1}^{N^{t}}$, where $\mathbf{x}^{t}_{i}\in\mathbb{R}^{H\times W\times C}$ is the $i$th input sample, $y^{t}_{i}\in\mathcal{C}^{t}$ is its desired prediction, and $N^{t}$ is the number of examples in task $t$. At each time step $t$, the model may only access $\mathcal{D}^{t}$ and is prohibited from storing or replaying data from any $\mathcal{D}^{u}$ with $u<t$. This constraint makes it challenging to avoid catastrophic forgetting, which updates on new tasks overwrite knowledge acquired on earlier ones.
Low-rank adaptation. LoRA [32] parameterizes a low-rank update to a pre-trained weight matrix $\mathbf{W}_{0}\in\mathbb{R}^{d_{out}\times d_{in}}$ by introducing two factors $\mathbf{B}\in\mathbb{R}^{d_{out}\times r}$ and $\mathbf{A}\in\mathbb{R}^{r\times d_{in}}$, such that $\Delta\mathbf{W}=\mathbf{B}\mathbf{A}$, where $r\ll\min(d_{in},d_{out})$. The updated weight matrix is then defined as:
@@
Mixture-of-Experts LoRA. Building on the Mixture-of-Experts (MoE) paradigm, a common framework of MoE-LoRA [85, 72] treats each LoRA as an independent expert and incrementally adds a new LoRA expert for each incoming task. After $T$ tasks, we have $T$ LoRA experts $\{(\mathbf{A}^{1},\mathbf{B}^{1}),\dots,(\mathbf{A}^{T},\mathbf{B}^{T})\}$. For an input token $x\in\mathbb{R}^{d_{\rm in}}$, the overall LoRA update in this framework is given by $\Delta\mathbf{W}=\sum\nolimits_{i=1}^{T}R(x)_{i}\mathbf{B}^{i}\,\mathbf{A}^{i},$ where the mixture weight $R(x)\in\mathbb{R}^{T}$ is produced by a learned router $R(\cdot)=\mathrm{softmax}(x\mathbf{W}_{r})$ and $\mathbf{W}_{r}\in\mathbb{R}^{d_{\rm in}\times T}$ contains the router’s trainable parameters. Each LoRA’s contribution is weighted by the learnable router, enabling the model to dynamically select and combine the most relevant low-rank updates for each token. In practice, a top-k masking is typically applied to mixture weights to enforce sparsity, activating only the $k$ most relevant experts and controlling computational overhead.

### v2

@@ Abstract
Continual learning (CL) with large pre-trained models is challenged by catastrophic forgetting and task interference. Existing LoRA-based Mixture-of-Experts (MoE) approaches mitigate forgetting by assigning and freezing task-specific adapters, but suffer from interference, redundancy, and ambiguous routing due to coarse adapter-level selection. However, this design introduces three key challenges: 1) Interference: Activating full LoRA experts per input leads to subspace interference and prevents selective reuse of useful components across tasks. 2) Redundancy: Newly added experts often duplicate or contradict existing knowledge due to unnecessary activation of unrelated ranks and insufficient reuse of relevant ones. 3) Ambiguity: Overlapping features across tasks confuse the router, resulting in unstable expert assignments. As more experts accumulate, earlier task routing degrades, accelerating forgetting. We propose MoRA, a Mixture-of-Rank Adaptive learning approaches with self-activated and sparse rank activation for CL. Unlike mixing multiple low-rank matrices, MoRA decomposes each rank-r update into r rank-one components, each treated as an independent expert, enabling fine-grained rank-one expert utilization while mitigating interference and redundancy. To avoid ambiguous routing, we propose that each rank-one expert can infer its own relevance via intermediate activations. Coupled with our proposed rank pruning and activation budgets, MoRA adaptively selects a sparse mixture of ranks per input. We validate MoRA on continual learning benchmarks using CLIP and language models, analyzing both in-domain learning and out-of-domain forgetting/generalization during fine-tuning. MoRA shows significant effectiveness in enhancing CL with PTMs, and improving generalization while mitigating forgetting.
@@ 1 Introduction
Continual learning (CL) (Hadsell et al., 2020; De Lange et al., 2021; Ding et al., 2022) aims to enable models to incrementally and efficiently acquire new knowledge from a stream of tasks, without catastrophic forgetting (Nguyen et al., 2019; McCloskey & Cohen, 1989) or the need for repeated fine-tuning on all previously seen data (Wang et al., 2022e; 2024). The emergence of large pre-trained models (PTMs), including Vision Transformer (ViT) for vision (Dosovitskiy et al., 2020), Language Model (LM) (Grattafiori et al., 2024; Raffel et al., 2020), and CLIP for vision–language embeddings (Radford et al., 2021), has spurred a wave exploring CL with PTMs (Xu et al., 2024; Wang et al., 2023). Despite the success of PTMs, CL on them across sequential tasks still faces key challenges, including preventing catastrophic forgetting and preserving, or even enhancing, the pre-trained capabilities to generalize to unseen or future cases (Qiao & Mahdavi, 2024).
@@
A common way to integrate new knowledge into PTMs is full fine-tuning (Zheng et al., 2023; Garg et al., 2023) or PEFT with a small set of learnable parameters (Yang et al., 2024b), such as LoRA (Hu et al., 2021) (Fig. 1(a)). However, continually updating the same parameters often leads to interference and forgetting, even with a single new task (Biderman et al., 2024). Leveraging LoRA’s regularization and flexibility, many CL methods adopt LoRA or its variants with additional strategies to mitigate forgetting (Yu et al., 2024; Wang et al., 2023; Qiao & Mahdavi, 2024).
@@
Inspired by mixture-of-experts (MoE) architectures (Shazeer et al., 2017; Lepikhin et al., 2020; Fedus et al., 2022; Dai et al., 2024) and the modular nature of LoRA adapters, MoE frameworks using LoRA adapters as experts have been studied (Dou et al., 2023; Wu et al., 2024b) and applied in CL (Fig. 1(b)) have been widely adopted in CL (Yu et al., 2024; Wang et al., 2024; Yang et al., 2024a; Chen et al., 2024; Li et al., 2025). These works either predefine an MoE with LoRA for CL (Yang et al., 2024a), or incrementally add experts (Wang et al., 2024) or task-specific routers (Yu et al., 2024), assuming MoE benefits CL by isolating task interference. Such methods (Wang et al., 2024; Qiao & Mahdavi, 2024; Rusu et al., 2016; Yu et al., 2024) freeze old components and add new ones (e.g., experts or routers) to reduce forgetting. Despite design differences, we collectively refer a plain and general design with LoRA-based MoE as MoE-LoRA. In MoE-LoRA models, each expert is a LoRA adapter with multiple ranks (subspaces), and the router selects among experts with each LoRA adapter as a unit. Despite their success, we argue that the coarse granularity of experts and routing introduces three key challenges that limit effective utilization:
(1) Interference across experts and limited knowledge reuse. A coarse-grained multi-rank LoRA expert is packed with a rich set of knowledge that must be fully activated or not at all. For each sample (i.e., token), any two experts may hold both complementary and conflicting knowledge, creating a dilemma: selecting both causes interference, while only selecting either one prevents reuse.
@@
(2) Learning redundancy and conflicts. As a result of (1), (newly added) experts often have to carry conflicting or redundant information, either from activating other experts with both useful and unnecessary content or from insufficient reuse of existing ones. This coarse selection exacerbates interference and forgetting during inference.
(3) Routing ambiguity and forgetting. With coarse-grained experts and routing, redundancy and conflicts hinder precise input–module assignment, leading to suboptimal mixtures (Lepikhin et al., 2020; Fedus et al., 2022; Dai et al., 2024). As more experts accumulate in CL, router assignments for earlier tasks grow unreliable, increasing forgetting.
The three challenges are intertwined causes and effects of coarse-grained experts and routing, limiting the potential of promising MoE-LoRA design.
@@
We propose MoRA, a Mixture-of-Rank Adaptive learning approaches with self-activated and sparsely selected ranks across the LoRAs added at different tasks. To address the limitations of coarse-grained MoE-LoRA methods, MoRA operates at a fine-grained level by treating each rank-one subspace within a LoRA adapter as an individual expert (or treating each rank-one LoRA as an expert). Expert selection and routing occur at the rank level, avoiding interference from whole LoRA blocks. This decomposition promotes specialization on narrower data regions, reduces redundancy, and enables effective knowledge reuse across tasks.
We introduce a self-activation mechanism to select a sparse set of ranks without relying on a separately trained router. By defining each rank-one component as an expert, we avoid a separate router for expert assignment: the rank-one $\mathbf{A}$ in LoRA can act as a weighting router for its $\mathbf{B}$. Each rank evaluates its relevance to the input and determines its contribution, grounded in the view of low-rank updates as linear key–value memory. This design unifies rank-one LoRA with fine-grained MoE while preserving the original characteristics, saving parameters (for additional router), reducing routing confusion, and further mitigating forgetting (Fig. 1(c)). Sparsity is enforced via temperature-scaled softmax together with top-k masking and rank pruning under a rank activation budget. As a complement, test-time thresholding further improves adaptability across cases. Leveraging the connection between PTMs and low-dimensional manifolds (Li et al., 2018; Aghajanyan et al., 2020), MoRA enables precise, input-specific adaptation through selective subspace updates. With fine-grained rank-level experts and sparse activation, MoRA enables continual learning by reusing old ranks and selectively activating a few new ones per sample, allowing efficient “little-by-little” adaptation with minimal redundancy. This sample-specific activation at inference also preserves pre-trained knowledge and improves generalization to unseen tasks.
We validate MoRA on CL with both vision–language CLIP and LMs, analyzing its impact on pre-trained knowledge during low-rank fine-tuning. MoRA delivers strong downstream performance with far fewer active parameters, greatly reduces forgetting of prior tasks and pre-trained knowledge, and even improves generalization to unseen domains.
@@ 2 Related Work
Continual learning enables sequential knowledge acquisition without forgetting. Experience replay (ER) methods (Luo et al., 2023; Aljundi et al., 2019b; Chaudhry et al., 2018a; Liu et al., 2020; Chaudhry et al., 2018b; Yan et al., 2022; 2021) interleave past examples with new data. Parameter regularization (Kirkpatrick et al., 2017; Aljundi et al., 2018; Zenke et al., 2017; Aljundi et al., 2019a; Jha et al., 2023) penalizes updates to critical weights. Dynamic networks (Zhou et al., 2024; Wang et al., 2022a; b; Zhou et al., 2022; Wang et al., 2024; McDonnell et al., 2024; Liang & Li, 2024; Wang et al., 2022e; d; Smith et al., 2023; Wang et al., 2022c; 2024) allocate new capacity on the fly and preserve dedicated pathways for prior tasks.
Continual learning of PTMs. For CL on vision–language CLIP model (Jha et al., 2024; Zhang et al., 2024b; Garg et al., 2023), methods like ZSCL (Zheng et al., 2023) retain zero-shot performance during adaptation, and follow-up work (Yu et al., 2024; Tang et al., 2025; Xu et al., 2024; Lu et al., 2024) continually fine-tunes while leveraging frozen pre-trained predictions. The X-TAIL benchmark (Xu et al., 2024) further challenges models by mixing domain labels at test time. In language models (LMs) (de Masson D’Autume et al., 2019; Qin & Joty, 2021; Razdaibiedina et al., 2023; Wang et al., 2023; Qiao & Mahdavi, 2024), continual learning uses capacity expansion or task-specific submodules to reduce interference.
Low‐rank adaptation (LoRA) (Hu et al., 2021) is widely used for parameter‐efficient fine‐tuning of large pre‐trained models. Building on this foundation, recent methods reformulate LoRA’s updates via SVD‐based initialization and dynamic rank scheduling (Zhang et al., 2024a; Meng et al., 2024; Ding et al., 2023; Zhang et al., 2023; Liu et al., 2024; Wu et al., 2024a), showing that task adaptation primarily lies in a compact subspace. In this work, we offer a complementary perspective by viewing both pre‐trained weight matrix and its low‐rank updates as a linear associative memory (Li et al., 2018; Aghajanyan et al., 2020; Kohonen, 1972; Anderson, 1972). Under this lens, a rank-$r$ update corresponds to writing $r$ new key–value entries into the pre-trained memory matrix, with each rank-one component acting as an independent memory slot.
Mixture‐of‐experts (MoE) in LoRA fine‐tuning. MoE increases model capacity by replacing the Transformer’s feed-forward layer with expert subnetworks, routing each input to a small subset and using load‐balancing losses to even out usage (Shazeer et al., 2017; Lepikhin et al., 2020; Fedus et al., 2022; Dai et al., 2024). This paradigm has been adapted for standard fine-tuning (Chen et al., 2023; Li et al., 2024; Dou et al., 2023) and CL (Yu et al., 2024; Wang et al., 2024; Yang et al., 2024a; Chen et al., 2024) with LoRA, treating each low‐rank adapter as an expert: new adapters are trained and frozen per task to prevent forgetting. In contrast, we decompose each rank‐$r$ update into $r$ rank‐11 components and compute an input‐dependent mixture over these fine‐grained experts, greatly enhancing both expert specialization and mixture diversity.
@@ 3.1 Preliminaries
Continual learning. In CL, a model sequentially learns $T$ tasks. For task $t\in{1,\dots,T}$, let $\mathcal{D}^{t}={(\mathbf{x}^{t}_{i},y^{t}_{i})}_{i=1}^{N^{t}}$, where $\mathbf{x}^{t}_{i}\in\mathbb{R}^{n\times d}$, $y^{t}_{i}\in\mathcal{C}^{t}$, and $N^{t}$ is the number of examples. In the memory-free setting, the model may access only $\mathcal{D}^{t}$ and cannot access data from any $\mathcal{D}^{u}$ with $u<t$.
Low-rank adaptation. LoRA (Hu et al., 2021) parameterizes a low‐rank update to a pre‐trained weight matrix $\mathbf{W}_{0}\in\mathbb{R}^{d_{out}\times d_{in}}$ by introducing two factors $\mathbf{B}\in\mathbb{R}^{d_{out}\times r}$ and $\mathbf{A}\in\mathbb{R}^{r\times d_{in}}$, such that $\Delta\mathbf{W}=\mathbf{B}\mathbf{A}$, where $r\ll\min(d_{in},d_{out})$. The updated weight matrix is then defined as:
@@
Mixture‐of‐Experts LoRA. Building on the Mixture‐of‐Experts (MoE) paradigm, a common generic framework of MoE‐LoRA (Yu et al., 2024; Wang et al., 2024) treats each LoRA as an independent expert. Suppose after $T$ tasks, we have $T$ LoRA experts $\{(\mathbf{A}^{1},\mathbf{B}^{1}),\dots,(\mathbf{A}^{T},\mathbf{B}^{T})\}$. For an input token $x\in\mathbb{R}^{d_{\rm in}}$, the overall LoRA update in this framework is given by $\Delta\mathbf{W}=\sum\nolimits_{i=1}^{T}R(x)_{i}\mathbf{B}^{i}\,\mathbf{A}^{i},$ where the mixture weight $R(x)\in\mathbb{R}^{T}$ is produced by a learned router $R(\cdot)=\mathrm{softmax}(x\mathbf{W}_{r})$ and $\mathbf{W}_{r}\in\mathbb{R}^{d_{\rm in}\times T}$ contains the router’s trainable parameters. Each LoRA’s contribution is weighted by the learnable router, enabling the model to dynamically select and combine the most relevant low‐rank updates for each token. In practice, a Top-$k$ masking is typically applied to mixture weights to enforce sparsity, activating only the $k$ most relevant experts and controlling computational overhead.
3.2 Model Weights and Updates as Linear Associative Memory Model
Weight matrix as linear associative memory. Large pre‐trained models (PTMs) have been shown to inhabit a low “intrinsic‐dimension” manifold compared to their full parameter count (Li et al., 2018; Aghajanyan et al., 2020). Any weight matrix of a pre-trained model $\mathbf{W}_{0}\in\mathbb{R}^{d_{\text{out}}\times d_{\text{in}}}$ can be treated as a linear associative memory model with a key-value structure (Kohonen, 1972; Anderson, 1972; Bau et al., 2020; Meng et al., 2022). Concretely, for any matrix $\mathbf{W}\in\mathbb{R}^{d_{\mathrm{out}}\times d_{\mathrm{in}}}$ and input $\mathbf{x}\in\mathbb{R}^{d_{\mathrm{in}}}$, the product $\mathbf{W}\,\mathbf{x}$ implements a content‐addressable memory read over $m$ stored key–value pairs $\{(\mathbf{K}_{i,:},\mathbf{V}_{:,i})\}_{i=1}^{m}$, with $m$ potentially large and varied across each weight matrix. Here, ${\mathbf{K}}=\begin{bmatrix}{\mathbf{K}}_{1,:}^{\top}&\cdots&{\mathbf{K}}_{m,:}^{\top}\end{bmatrix}^{\top}\in\mathbb{R}^{m\times d_{\text{in}}}$ and $\mathbf{V}$$\begin{bmatrix}{\mathbf{V}}_{:,1}&\cdots&{\mathbf{V}}_{:,m}\end{bmatrix}\in\mathbb{R}^{d_{\text{out}}\times m}$ with $\mathbf{K}_{i,:}\in\mathbb{R}^{1\times d_{\text{in}}}$ is the $i$-th row of $\mathbf{K}$} and ${\mathbf{V}}_{:,i}\in\mathbb{R}^{d_{\text{out}}\times 1}$ is the $i$-th column of $\mathbf{V}$}. Given any input $\mathbf{x}$, the output $\mathbf{y}\in\mathbb{R}^{d_{\text{out}}}$ is computed as
$\mathbf{y}=\mathbf{W}\mathbf{x}\approx\sum\nolimits_{i=1}^{m}\mathbf{V}_{:,i}(\mathbf{K}_{i,:}\mathbf{x}).$ (2)
In this formulation, the forward pass through the weight matrix with the input decomposes into two steps: (1) computing response scores or assignment weights via the key vectors $\mathbf{K}_{i,:}$, and (2) aggregating the corresponding value vectors $\mathbf{V}_{:,i}$ according to those weights.